OpenStudy (anonymous):

A tumor is a collection of cancerous cells. Suppose each cancerous cell has a radius of 5 times 10^{-3} cm, and that the doubling time for a population of cancerous cells is 50 hours. If the tumor has an initial volume of 0.1 cm^3, how many cancerous cells does it contain? (For simplicity, assume that the tumor is entirely made of cells, that there is no space between them.) Round your answer up to the next whole number. Find a formula for V(t), the total volume of the cancerous cells, after t hours. V(t)= How many hours will it take before the volume of the tumor reaches 1 cm^

OpenStudy (anonymous):

I am stuck on the formula for V(t)

OpenStudy (anonymous):

what is the initial amount?

OpenStudy (anonymous):

what ever it is, since it doubles in 50 hours you can use \[A_0\times 2^{\frac{t}{50}}\]

OpenStudy (anonymous):

I had some help on this before and got this far with it V(t) = Voe^kt 4/3 * pi * r^3 = 5.293598 * 10^-7 1.91 * 10^5 = 191,000

OpenStudy (anonymous):

Initial amount is 191,000 with an initial volume of 0.1cm^3

OpenStudy (anonymous):

I need to find out what the formula for V(t) is and How many hours will it take before the volume of the tumor reaches 1 cm^3? Round your answer up to the next whole numbe